# Community detection in heterogeneous signed networks

### v1

@@ 4 Theory
In this section, we establish the asymptotic properties of the proposed SBBM in terms of both probability estimation and community detection. Denote the true parameters as $(\bm{\beta}^{*+},\bm{\beta}^{*-},\bm{\gamma}^{*+},\bm{\gamma}^{*-},\bm{Z}^{*})\in\mathcal{P}$ with $\bm{\beta}^{*+}>\bm{\beta}^{*-}$ and $\bm{\gamma}^{*+}<\bm{\gamma}^{*-}$, and $\bm{\Theta}^{*}=(\bm{\Theta}^{*+},\bm{\Theta}^{*-})$ correspondingly.
4.1 Consistency in probability estimation
For each $i,j\in[n]$, denote the negative log-likelihood of $a_{ij}$ as
$L(\theta_{ij};a_{ij})=-(a_{ij}^{+}\theta_{ij}^{+}+a_{ij}^{-}\theta_{ij}^{-})+\log\big(1+e^{\theta_{ij}^{+}}+e^{\theta_{ij}^{-}}\big),$
where $\theta_{ij}=(\theta_{ij}^{+},\theta_{ij}^{-})$, and $L(\bm{\Theta};\bm{A})=\frac{2}{n(n+1)}\sum_{i\leq j}L(\theta_{ij};a_{ij})$. Let $e_{L}(\bm{\Theta},\bm{\Theta}^{*})=\mathbb{E}\big(L(\bm{\Theta};\bm{A})-L(\bm{\Theta}^{*};\bm{A})\big)$, which serves as an intermediate error metric measuring the accuracy of $\bm{\Theta}$ in estimating $\bm{\Theta}^{*}$.
Assumption 1
There exist constants $M_{1}<M_{2}$ such that both $\theta_{ij}^{+}$ and $\theta_{ij}^{-}$ are contained in $M_{1},M_{2}]$ for each $i,j\in[n]$.
@@ 4.1 Consistency in probability estimation
Proposition 3
For any $\bm{\Theta}$ that satisfies Assumption 1, it holds that
$\frac{2}{n(n+1)}\sum_{i\leq j}\mathbb{E}[L(\theta_{ij},a_{ij})-L(\theta^{*}_{ij},a_{ij})]^{2}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}e_{L}(\bm{\Theta},\bm{\Theta}^{*}),$
$\frac{1}{n^{2}}\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ $\leq 48e^{\max\{|M_{1}|,|M_{2}|\}}e_{L}(\bm{\Theta},\bm{\Theta}^{*}).$
Proposition 3 paves a bridge between the first and second moments of $L(\bm{\Theta},\bm{A})-L(\bm{\Theta}^{*},\bm{A})$, which is crucial for establishing the weak consistency of $\widehat{\bm{\Theta}}$ in terms of $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$. Furthermore, it also relates $\big\|{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}$ to $e_{L}(\bm{\Theta},\bm{\Theta}^{*})$, leading to the strong consistency of $\widehat{\bm{\Theta}}$ established in the following Theorem 2.
Theorem 2
[…]
$\ P\Big(\frac{1}{n^{2}}\big\|\widehat{\bm{\Theta}}-\bm{\Theta}^{*}\big\|^{2}_{F}\geq 48e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\Big)\leq 2\exp(-c_{1}Kn\log n),$
when $n$ is large enough, provided that $\epsilon_{n}^{2}\simeq\frac{Ke^{\max\{|M_{1}|,|M_{2}|\}}\min\{2e^{M_{2}}+e^{2M_{2}},1\}\log n}{n}\prec 1$.
Theorem 2 guarantees that $\bm{\Theta}^{*}$ can be consistently estimated if $K\simeq 1$ and $\min\{2e^{M_{2}}+e^{2M_{2}},1\}e^{2\max\{|M_{1}|,|M_{2}|\}}n^{-1}\log n$ is vanishing. In particular, if $M_{2}-M_{1}\preceq 1$ and both of them diverge to -$\infty$, then it is required that $\min\{p_{ij}^{*+},p_{ij}^{*-}\}\succ\frac{\log n}{n}$ for each $i,j\in[n]$, which matches the best sparsity requirement in the literature [lei2015consistency; ma2020universal; zhang2022directed]. More importantly, when $M_{2}-M_{1}\succ 1$, $p_{ij}^{*+}$ and $p_{ij}^{*-}$ can behave differently to accommodate edge heterogeneity. Specifically, letting $M_{1}\to$-$\infty$ and $M_{2}\to\infty$, Theorem 2 still holds when $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$, resulting in a heterogeneous case with $p_{ij}^{*+},p_{ij}^{*-}\in(n^{-1/2}(\log n)^{1/2},1)$ for each $i,j\in[n]$. This is consistent with the heterogeneous result for the latent space model [ma2020universal; zhang2022joint], and better than that for the classical $\beta$-model [shao20212], where the connecting probabilities can only vary in $(n^{-1/4},1)$.
@@ 4.2 Consistency in community detection
We next turn to establish the asymptotic properties of $\widehat{\bm{Z}}$, whose community detection error is defined as
$d(\widehat{\bm{Z}},\bm{Z}^{*})=\min_{\bm{\Pi}}\frac{1}{n}\|\widehat{\bm{Z}}\bm{\Pi}-\bm{Z}^{*}\|_{F}^{2},$ (4)
@@
where minimization is taken over all permutation matrices $\bm{\Pi}\in\{0,1\}^{K\times K}$. The quantity $d(\widehat{\bm{Z}},\bm{Z}^{*})$ measures the minimal fraction of misclassified nodes, up to permutation, and is commonly used in the community detection literature [lei2015consistency].
Assumption 2
There exists a constant $M_{\gamma}$ such that $\|\bm{\gamma}\|_{\infty}\leq M_{\gamma}$. Let $\mathcal{S}_{i}^{*+}=\big\{j\in[n]:\ |\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}\big\}$ and $\mathcal{S}_{i}^{*-}=\big\{j\in[n]:\ |\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}\big\}$ for some $c_{\eta}>0$, then $\max_{i\in[n]}|\mathcal{S}_{i}^{*+}|\prec\frac{n}{K}$ or $\max_{i\in[n]}|\mathcal{S}_{i}^{*-}|\prec\frac{n}{K}$.
Assumption 2 assures that the number of pairs of $(i,j)$ satisfying $|\beta_{i}^{*+}+\beta_{j}^{*+}|<2M_{\gamma}+c_{\eta}$ or $|\beta_{i}^{*-}+\beta_{j}^{*-}|<2M_{\gamma}+c_{\eta}$ is substantially less than $n/K$. Note that $|\beta_{i}^{*+}+\beta_{j}^{*+}|\geq 2M_{\gamma}+c_{\eta}$ and $|\beta_{i}^{*-}+\beta_{j}^{*-}|\geq 2M_{\gamma}+c_{\eta}$ imply that $\min\{|\eta_{i}^{*+}+\eta_{j}^{*+}|,|\eta_{i}^{*-}+\eta_{j}^{*-}|\}\geq c_{\eta}$, thus $c_{\eta}$ can be regarded as a measurement of the difference in the level of connection probability within the same community and those across different communities. For any $(\bm{\gamma}^{+},\bm{\gamma}^{-},\bm{\eta}^{+},,\bm{\eta}^{-},\bm{Z})\in\mathcal{P}$, let $\mathcal{N}_{k}=\{i\in[n]:z_{ik}=1\}$ and $n_{k}=|\mathcal{N}_{k}|$ for each $k\in[K]$.
Proposition 4
Suppose all assumptions in Theorem 2 and Assumption 2 hold, and that $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$. Then, with probability at least $1-2\exp(-c_{1}Kn\log n)$, there does not exist $k_{1}\neq k_{2}\in[K]$ and $k\in[K]$ such that $|\mathcal{N}_{k_{1}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq|\mathcal{N}_{k_{2}}^{*}\cap\widehat{\mathcal{N}}_{k}|\simeq\frac{n}{K}$.
The requirement of $K^{2}c_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\prec 1$ implies that $c_{\eta}$ can not be too small, which can be interpreted as a minimal signal strength condition. Proposition 4 assures that any estimated community $\widehat{\mathcal{N}}_{k}$ does not substantially overlap with more than one true community with high probability, which further leads to the estimation consistency of $\widehat{\bm{Z}}$ as shown in Theorem 3.
Theorem 3
@@
Suppose all assumptions in Proposition 4 hold, then there exists a positive constant $c_{2}$ such that when $n$ is sufficiently large,
$P\big(d(\widehat{\bm{Z}},\bm{Z}^{*})\geq c_{2}Kc_{\eta}^{-2}e^{\max\{|M_{1}|,|M_{2}|\}}\epsilon_{n}^{2}\big)\leq 2\exp(-c_{1}Kn\log n).$
Theorem 3 establishes the convergence rate of $d(\widehat{\bm{Z}},\bm{Z}^{*})$ for community detection. Particularly, in the homogeneous case with $\max\{|M_{1}|,|M_{2}|,K,c_{\eta}\}\simeq 1$, the convergence rate becomes $n^{-1}\log n$, which matches with the best community detection result in literature [lei2015consistency; zhang2022directed]. In the heterogeneous case with $e^{M_{1}}\succ\sqrt{\frac{\log n}{n}}$ and $e^{M_{2}}\prec\sqrt{\frac{n}{\log n}}$ where $p_{ij}^{*}$ is allowed to vary in $(n^{-1/2}(\log n)^{1/2},1)$ for each $i\leq j$, $c_{\eta}$ can diverge to infinity at a rate no faster than $\log n$ and Theorem 3 continues to hold with $d(\widehat{\bm{Z}},\bm{Z}^{*})=O_{p}\big((\log n)^{-2}\big)$.
@@ 5 Numerical experiments
In this section, we examine the numerical performance of the proposed SBBM model, and compare it against three popular competitors in the literature, including the signed Laplacian-based method [kunegis2010spectral], the SPONGE algorithm based on signed SBM [cucuringu2021regularized], and the joint estimation method based on the joint inner product model [tang2024population], denoted as SLP, SPONGE and JIM, respectively. All methods are evaluated in terms of community detection accuracy, while JIM and SBBM are further compared in terms of probability estimation.
Note that the community detection error in (4) requires enumeration of all possible permutations, which is computationally infeasible. We thus employ the clustering error in wang2010consistent to evaluate the community detection accuracy. Particularly, let $\bm{\psi}^{*}=(\psi_{1}^{*},\dots,\psi_{n}^{*})$ and $\widehat{\bm{\psi}}=(\widehat{\psi}_{1},\dots,\widehat{\psi}_{n})$ denote the true and estimated community assignments, then the community detection error of $\widehat{\bm{\psi}}$ can be evaluated as
$\text{Err}(\widehat{\bm{\psi}},\bm{\psi}^{*})=\frac{2}{n(n-1)}\sum_{i<j}\mathbb{I}\big(\mathbb{I}(\psi_{i}^{*}=\psi_{j}^{*})+\mathbb{I}(\widehat{\psi}_{i}=\widehat{\psi}_{j})=1\big).$
Further, we define the estimation error of the probability matrix $\bm{P}^{*+}=(p^{*+}_{ij})_{i,j=1}^{n}$ and $\bm{P}^{*-}=(p^{*-}_{ij})_{i,j=1}^{n}$ as
[…]
@@ 5.1 Simulated examples
Example 1. Following the data generating scheme in tang2024population, we first generate $\bm{X}\in\mathbb{R}^{n\times 2}$ with rows independently set to $(1,0)$ or $(0,1)$ with equal probability $0.5$, and then set $\bm{X}^{*}=(\bm{I}_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\top})\bm{X}$. For each node $i\in[n]$, we generate $\alpha_{i}^{*}\sim\mathrm{Unif}(1,3)$ and let $v_{i}^{*}=(1,-1)^{\top}\bm{x}_{i}^{*}/\sqrt{2}$, where $\bm{x}_{i}^{*}$ denotes the $i$-th row of $\bm{X}^{*}$. Edge probabilities are then specified by $\Pr(|A_{ij}|=1)=\sigma(\mu^{*}+\alpha_{i}^{*}+\alpha_{j}^{*}-2\bar{\alpha}^{*}+\bm{x}_{i}^{*\top}\bm{x}_{j}^{*})$ and $\Pr(A_{ij}=1\mid|A_{ij}|=1)=\sigma(v_{i}^{*}v_{j}^{*})$, where $\bar{\alpha}^{*}=n^{-1}\sum_{i=1}^{n}\alpha_{i}^{*}$ and $\sigma(x)=(1+e^{-x})^{-1}$. Varying $\mu^{*}\in\{-3.5,-3.0,-2.5\}$ produces strong-balanced signed networks with different sparsity.
Example 2. In this example, for each $i\in[n]$, we first generate $\beta_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, $\gamma_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$, where $\delta_{i}\sim U(0,2)$. Then we vary $K\in\{4,6,8\}$ and randomly assign each node to one of the $K$ communities uniformly. The matrices $\bm{\Theta}^{*}$ and $\bm{P}^{*}$ are constructed as in (1) and the signed networks are generated accordingly, resulting in weak-balanced signed networks with different numbers of communities.
Example 3. In this example, we set $K=4$. For each $i\in[n]$, we generate $\beta_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$, $\gamma_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$ with $\sigma^{2}\in\{1,2,3\}$. We then generate $\delta_{i}\sim U(0,2)$ and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$. The signed networks are subsequently generated as in Example 2, resulting in a weak-balanced network with various node heterogeneity.
We consider $n\in\{500,1000\}$ in all examples. The averaged probability estimation errors and community detection errors, along with their standard errors, over 50 independent replications are reported in Tables 1-6.
@@
500 -3.5 0.0910(0.0122) 0.4172(0.0069) 0.3950(0.0106) 0.2073(0.0242)
-3.0 0.0271(0.0014) 0.2822(0.0059) 0.2164(0.0049) 0.1060(0.0189)
-2.5 0.0078(0.0007) 0.1715(0.0050) 0.1169(0.0034) 0.0361(0.0106)
1000 -3.5 0.0133(0.0007) 0.2016(0.0029) 0.1515(0.0029) 0.0723(0.0162)
-3.0 0.0025(0.0003) 0.1070(0.0023) 0.0711(0.0018) 0.0230(0.0058)
-2.5 0.0004(0.0001) 0.0440(0.0012) 0.0257(0.0009) 0.0161(0.0100)
Table 2: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 1. The better performer in each scenario is boldfaced.
$n$ $\mu$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 -3.5 0.6085(0.1521) 0.8727(0.2557) 0.5546(0.0126) 0.4829(0.0106)
-3.0 0.3136(0.0014) 0.3964(0.0017) 0.4163(0.0091) 0.3848(0.0094)
-2.5 0.2582(0.0012) 0.3199(0.0011) 0.3187(0.0067) 0.3033(0.0079)
1000 -3.5 0.2789(0.0008) 0.3527(0.0010) 0.3726(0.0063) 0.3528(0.0069)
-3.0 0.2277(0.0007) 0.2850(0.0007) 0.2819(0.0032) 0.2805(0.0067)
-2.5 0.1888(0.0005) 0.2350(0.0005) 0.2170(0.0048) 0.2140(0.0065)
@@
500 4 0.0310(0.0041) 0.3201(0.0063) 0.2276(0.0073) -
6 0.0425(0.0036) 0.3621(0.0049) 0.2692(0.0050) -
8 0.0508(0.0039) 0.3800(0.0056) 0.2886(0.0051) -
1000 4 0.0091(0.0006) 0.3128(0.0050) 0.2141(0.0060) -
6 0.0123(0.0011) 0.3497(0.0037) 0.2600(0.0038) -
8 0.0152(0.0012) 0.3632(0.0031) 0.2736(0.0032) -
Table 4: Averaged relative probability estimation errors and their standard errors of SBBM (left) and JIM (right) over 50 independent replications in Example 2. The better performer in each scenario is boldfaced.
$n$ $K$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$ $\text{Err}(\widehat{\bm{P}}^{+},\bm{P}^{*+})$ $\text{Err}(\widehat{\bm{P}}^{-},\bm{P}^{*-})$
500 4 0.2265(0.0083) 0.1945(0.0049) 1.0216(0.0093) 0.8595(0.0035)
6 0.2841(0.0119) 0.1826(0.0038) 1.1016(0.0123) 0.8305(0.0033)
8 0.3409(0.0128) 0.1819(0.0044) 1.1649(0.0149) 0.8164(0.0033)
1000 4 0.1546(0.0035) 0.1344(0.0024) 1.0343(0.0096) 0.8587(0.0026)
6 0.1788(0.0049) 0.1225(0.0020) 1.1188(0.0130) 0.8290(0.0025)
8 0.1993(0.0053) 0.1168(0.0016) 1.1869(0.0159) 0.8151(0.0024)

### v2

@@ 4 Theory
In this section, we establish the asymptotic properties of the proposed SBBM in terms of the estimation error of $(\widehat{\bm{\Theta}}^{+},\widehat{\bm{\Theta}}^{-})$ obtained in (3) and the community assignment error of $\widehat{\sigma}$ obtained in Section 3.2.
4.1 Consistency in parameter estimation
We first introduce several auxiliary quantities to facilitate the theoretical analysis. The true parameters are denoted by $\bm{\beta}^{*+},\bm{\beta}^{*-},\bm{\gamma}^{*+},\bm{\gamma}^{*-}$ and $\sigma^{*}$ with $\bm{\beta}^{*+}>\bm{\beta}^{*-}$ and $\bm{\gamma}^{*+}<\bm{\gamma}^{*-}$ component-wisely. Define $\bm{\Theta}^{*}=(\bm{\Theta}^{*+},\bm{\Theta}^{*-})$ and let $\bm{\Theta}_{\Delta}^{*}=\bm{\Theta}^{*+}-\bm{\Theta}^{*-}$. Finally, let $p^{*+}_{ij}=e^{\theta_{ij}^{*+}}/(1+e^{\theta_{ij}^{*+}}+e^{\theta_{ij}^{*-}})$, $p^{*-}_{ij}=e^{\theta_{ij}^{*-}}/(1+e^{\theta_{ij}^{*+}}+e^{\theta_{ij}^{*-}})$ be the true edge probabilities and $\bm{P}^{*+}=(p^{*+}_{ij})_{i,j=1}^{n},\bm{P}^{*-}=(p^{*-}_{ij})_{i,j=1}^{n}$.
Assumption 1.
The parameter $\bm{\Theta}=(\bm{\Theta}^{+},\bm{\Theta}^{-})$ satisfies that $M_{1}\leq\bm{\Theta}^{+}_{ij},\bm{\Theta}^{-}_{ij}\leq-M_{2}$ for all $1\leq i,j\leq n$, where $M_{1}$ and $M_{2}$ are allowed to diverge with $n$.
@@ 4.1 Consistency in parameter estimation
Theorem 2.
Suppose Assumption 1 holds and $\min\{\lambda_{n}^{+},\lambda_{n}^{-}\}\geq c_{0}\sqrt{\max\{ne^{-M_{2}},\log n\}}$ for a large enough constant $c_{0}$, it then holds true with probability at least $1-n^{-c}$ that
$\frac{1}{n^{2}}\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}^{*}_{\Delta}\|^{2}_{F}\leq\frac{c^{\prime}K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{{\lambda_{n}^{+}}^{2},{\lambda_{n}^{-}}^{2}\}.$ (4)
for some positive constants $c$ and $c^{\prime}$. Furthermore, if we set $\lambda_{n}^{+},\lambda_{n}^{-}\simeq\sqrt{\max\{ne^{-M_{2}},\log n\}}$, then with the same probability, we have
$\frac{1}{n^{2}}\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}^{*}_{\Delta}\|^{2}_{F}\leq\frac{c^{\prime\prime}K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{ne^{-M_{2}},\log n\}.$ (5)
for some positive constant $c^{\prime\prime}$.
Theorem 2 guarantees that $\bm{\Theta}^{*}_{\Delta}$ can be consistently estimated if $\lambda^{+}$ and $\lambda^{-}$ are sufficiently large but $\frac{K}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{{\lambda_{n}^{+}}^{2},{\lambda_{n}^{-}}^{2}\}$ vanishes. Its optimal convergence rate is attained when $\lambda_{n}^{+},\lambda_{n}^{-}\simeq\sqrt{\max\{ne^{-M_{2}},\log n\}}$. If $K\simeq 1$ and $M_{1}-M_{2}\preceq 1$ but both diverge to $\infty$, then it is required that $\min\{p_{ij}^{*+},p_{ij}^{*-}\}\succ\frac{\sqrt{\log n}}{n}$ for each $i,j\in[n]$, which matches the best sparsity requirement in the literature [lei2015consistency; ma2020universal; zhang2022directed]. Furthermore, when $M_{1}-M_{2}\succ 1$, $p_{ij}^{*+}$ and $p_{ij}^{*-}$ can behave differently to accommodate edge heterogeneity. Specifically, letting $e^{-M_{1}}\prec 1$ and $\frac{\sqrt{\log n}}{n}\prec e^{-M_{2}}\preceq 1$, Theorem 2 still holds when $e^{-M_{1}}\succ n^{-1/2}$, resulting in a heterogeneous case with $p_{ij}^{*+},p_{ij}^{*-}\in(n^{-1/2},1)$ for each $i,j\in[n]$. This is consistent with the heterogeneous result for the latent space model [ma2020universal; zhang2022joint], and better than that for the classical $\beta$-model [shao20212], where the connecting probabilities can only vary in $(n^{-1/4},1)$.
@@ 4.2 Consistency in community detection
Let $K\simeq 1$ and $\frac{1}{n^{2}}e^{2M_{1}}(1+e^{-M_{2}})^{4}\max\{ne^{-M_{2}},\log n\}\prec 1$. We now turn to establish the asymptotic properties of $\widehat{\sigma}$. Let
$d(\widehat{\sigma},\sigma^{*})=\min_{\pi}\frac{1}{n}\sum_{i\in[n]}\big(\pi\circ\widehat{\sigma}(i)-\sigma^{*}(i)\big)^{2}$ (6)
@@
be the community detection error, where $\pi:[K]\rightarrow[K]$ is the permutation mapping. Then, $d(\widehat{\sigma},\sigma^{*})$ measures the minimal fraction of misclassified nodes up to permutation, and is commonly used in the literature [lei2015consistency].
Assumption 2.
Let $\bm{\eta}_{\Delta}^{*}=(\bm{\beta}^{*+}-\bm{\gamma}^{*+})-(\bm{\beta}^{*-}-\bm{\gamma}^{*-})=(\eta_{\Delta,i}^{*})_{i=1}^{n}$. For each $k\in[K]$, let $s_{k}^{*}=\sum_{\sigma^{*}(i)=k}1$, $t_{k}^{*}=\sum_{\sigma^{*}(i)=k}\eta_{\Delta,i}^{*}$ and $u_{k}^{*}=\sum_{\sigma^{*}(i)=k}(\eta_{\Delta,i}^{*})^{2}$, then there exist $\bar{s}_{k}>0$ and $\bar{t}_{k},\bar{u}_{k}\in\mathbb{R}$ and $\epsilon_{n}\prec 1$ such that $\max_{k}\{|s_{k}^{*}/n-\bar{s}_{k}|,|t_{k}^{*}/n-\bar{t}_{k}|,|u_{k}^{*}/n-\bar{u}_{k}|\}\leq\epsilon_{n}$. Moreover, elements in $\{\bar{t}_{k}\pm\sqrt{\bar{s}_{k}\bar{u}_{k}},k\in[K]\}$ are all different.
Assumption 2 imposes an empirical moments convergence condition. Such uniform stabilization assumptions are standard in classical large-sample theory [andersen1982cox]. The requirement on the convergence of $s_{k}^{*}/n$ is analogous to the balanced community assumption in lei2015consistency. Furthermore, Assumption 2 guarantees that lines characterized in Proposition 3 converge to distinct lines as $n$ diverges, see Lemma S5 in supplementary material for details.
Assumption 3.
For $\{\eta_{\Delta,i}^{*}\}$, $i=1,\dots,n$,
  • (i)
    There exists $v:(0,1]\rightarrow\mathbb{R}^{+}$ such that for any $\epsilon\in(0,1]$,
    $v(\epsilon)=\min_{k\in[K]}\liminf_{C\subset\{i,\sigma^{*}(i)=k\},|C|\geq\epsilon n}\sqrt{\frac{1}{|C|}\sum_{i\in C}(\eta_{\Delta,i}^{*}-\frac{1}{|C|}\sum_{i\in C}\eta_{\Delta,i}^{*})^{2}}>0;$
  • (ii)
    If $K=2$, there exists $M>0$, such that $\|\bm{\eta}_{\Delta}^{*}\|_{\infty}<M$.
Assumption 3 (i) is a uniform non-degeneracy condition, which rules out the pathological case where $\eta_{\Delta,i}^{*}$ is nearly constant on any subset of a community, which aligns with the requirement of the identifiability condition in Theorem 1. Assumption 3 (ii) requires a uniform upper bound for $|\eta_{\Delta,i}^{*}|$ when there are only two communities. Both assumptions can be satisfied if $\eta_{\Delta,i}^{*}$’s within the community come independently from some continuous distribution. For example, if $\eta_{\Delta,i}^{*}$ follows a uniform distribution on $[0,1]$, then $v(\epsilon)\simeq\epsilon$.
Theorem 3.
Let $\gamma_{n}$ be the smallest nonzero eigenvalue of $\bm{J}\bm{\Theta}^{*}_{\Delta}\bm{J}$. Suppose all assumptions in Theorem 2 and Assumptions 2-3 hold, then it holds true with probability $1-n^{-c}$ for positive constant $c$ that
  • (i)
@@
    When $K\geq 3$, then there exists a positive constant $C$ that depends on $\{\bar{s}_{k}\}_{k\in[K]}$, $\{\bar{t}_{k}\}_{k\in[K]}$, $\{\bar{u}_{k}\}_{k\in[K]}$ and $K$ such that for sufficiently large $n$,
    $d(\widehat{\sigma},\sigma^{*})\cdot v(d(\widehat{\sigma},\sigma^{*}))\leq C\sqrt{\frac{\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}_{\Delta}^{*}\|_{F}^{2}}{\gamma^{2}_{n}}+\epsilon_{n}};$ (7)
  • (ii)
    When $K=2$, then there exists a positive constant $C$ that depends only on $\{\bar{s}_{k}\}_{k\in[K]}$, $\{\bar{t}_{k}\}_{k\in[K]}$, $\{\bar{u}_{k}\}_{k\in[K]}$, $K$ and $M$ such that for sufficiently large $n$,
    $\frac{d^{3}(\widehat{\sigma},\sigma^{*})\cdot v^{2}(d(\widehat{\sigma},\sigma^{*}))}{d(\widehat{\sigma},\sigma^{*})+v^{2}(d(\widehat{\sigma},\sigma^{*}))}\leq C\left(\frac{\|\widehat{\bm{\Theta}}_{\Delta}-\bm{\Theta}_{\Delta}^{*}\|_{F}^{2}}{\gamma^{2}_{n}}+\epsilon_{n}\right).$ (8)
Theorem 3 shows that the community detection error depends on the estimation error of $\widehat{\bm{\Theta}}_{\Delta}$, the approximation error $\epsilon_{n}$ and the variability of $\bm{\eta}_{\Delta}^{*}$. By Lemma S5, we have $\gamma_{n}\simeq n$, thus the right-hand sides of (7) and (8) converge to zero, and it immediately follows that $d(\widehat{\sigma},\sigma^{*})=o_{p}(1)$ by noting that $v(\cdot)$ is non-decreasing.
@@ 5 Numerical experiments
In this section, we examine the numerical performance of the proposed SBBM model, and compare it against three popular competitors in the literature, including the signed Laplacian-based method [kunegis2010spectral], the SPONGE algorithm based on signed SBM [cucuringu2021regularized], and the joint estimation method based on the joint inner product model [tang2024population], denoted as SLP, SPONGE and JIM, respectively. All methods are evaluated in terms of the community detection error defined in (6).
@@ 5.1 Simulated examples
Example 1. Following the data generating scheme in tang2024population, we first generate $\bm{X}\in\mathbb{R}^{n\times 2}$ with rows independently set to $(1,0)$ or $(0,1)$ with equal probability $0.5$, and then set $\bm{X}^{*}=(\bm{I}_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\top})\bm{X}$. For each node $i\in[n]$, we generate $\alpha_{i}^{*}\sim\mathrm{Unif}(1,3)$ and let $v_{i}^{*}=(1,-1)^{\top}\bm{x}_{i}^{*}/\sqrt{2}$, where $\bm{x}_{i}^{*}$ denotes the $i$-th row of $\bm{X}^{*}$. Edge probabilities are then specified by $\Pr(|A_{ij}|=1)=\sigma(\mu^{*}+\alpha_{i}^{*}+\alpha_{j}^{*}-2\bar{\alpha}^{*}+\bm{x}_{i}^{*\top}\bm{x}_{j}^{*})$ and $\Pr(A_{ij}=1\mid|A_{ij}|=1)=\sigma(v_{i}^{*}v_{j}^{*})$, where $\bar{\alpha}^{*}=n^{-1}\sum_{i=1}^{n}\alpha_{i}^{*}$ and $\sigma(x)=(1+e^{-x})^{-1}$. Varying $\mu^{*}\in\{-4.0,-3.5,-3.0\}$ produces strong-balanced signed networks with different sparsity.
Example 2. In this example, for each $i\in[n]$, we first generate $\beta_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, $\gamma_{i}^{*-}\sim\mathcal{N}(-\frac{7}{2},9)$, and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$, where $\delta_{i}\sim U(0,2)$. Then we vary $K\in\{3,4,5\}$ and randomly assign each node to one of the $K$ communities uniformly. The matrices $\bm{\Theta}^{*}$ and $\bm{P}^{*}$ are constructed as in (1) and the signed networks are generated accordingly, resulting in weak-balanced signed networks with different numbers of communities.
Example 3. In this example, we set $K=4$. For each $i\in[n]$, we generate $\beta_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$, $\gamma_{i}^{*-}\sim\mathcal{N}(-3,\sigma^{2})$ with $\sigma^{2}\in\{2,3,4\}$. We then generate $\delta_{i}\sim U(0,2)$ and set $\beta_{i}^{*+}=\beta_{i}^{*-}+\delta_{i}$ and $\gamma_{i}^{*+}=\gamma_{i}^{*-}-\delta_{i}$. The signed networks are subsequently generated as in Example 2, resulting in a weak-balanced network with various node heterogeneity.
We consider $n\in\{500,1000\}$ in all examples. The averaged community detection errors, along with their standard errors, over 50 independent replications are reported in Tables 1-3.
@@
500 -4.0 0.1392(0.0076) 0.4481(0.0066) 0.4370(0.0080) 0.2618(0.0204)
-3.5 0.0909(0.0025) 0.3067(0.0095) 0.2992(0.0160) 0.1536(0.0233)
-3.0 0.0632(0.0022) 0.1712(0.0047) 0.1239(0.0033) 0.0692(0.0153)
1000 -4.0 0.0860(0.0024) 0.2357(0.0051) 0.2187(0.0137) 0.1332(0.0201)
-3.5 0.0578(0.0019) 0.1138(0.0019) 0.0826(0.0017) 0.0481(0.0140)
-3.0 0.0414(0.0011) 0.0567(0.0013) 0.0369(0.0010) 0.0121(0.0032)
@@
500 3 0.0379(0.0024) 0.0857(0.0109) 0.1296(0.0063) -
4 0.0687(0.0046) 0.3170(0.0063) 0.2302(0.0072) -
5 0.1215(0.0098) 0.3799(0.0057) 0.2938(0.0077) -
1000 3 0.0174(0.0009) 0.0363(0.0080) 0.0955(0.0062) -
4 0.0312(0.0014) 0.3114(0.0051) 0.2194(0.0080) -
5 0.0591(0.0049) 0.3750(0.0038) 0.2910(0.0056) -